OpenStudy (anonymous):

What is the equation of the line that passes through the points (–12, –8) and (–17, –16)?

OpenStudy (hyuna301):

y = mx + c m = (y2 - y1)/(x2 - x1) So here m = 8/5 y = 8/5 x + c

OpenStudy (anonymous):

answer choices y-17=-8/5(x-16) y-16=8/5(x-17) y+16=-4(x+10) y+6=8/5(x+17)

OpenStudy (hyuna301):

If two points are given then use eq ( y - y1) = m( x - x1) so, eq is y +16 = 8/5( x + 17)